OpenStudy (anonymous):

I am answering a problem and I get 9. My friend says the answer is 1. I believe I am right because I am using PEMDAS. The problem is: 6/2(1+2)=? This is, well, my friend is driving me insane. Please help.

OpenStudy (anonymous):

i get 1

OpenStudy (anonymous):

you follow pemdas so you do parenthesis first so 2 + 1 = 3 then you do multiplication 3(2) = 6 then division 6/6 = 1

OpenStudy (anonymous):

probably because your friend is answering this \[\frac{ 6 }{2 }(2+1)\] and you're answering this \[\frac{ 6 }{2(2+1)}\]

OpenStudy (anonymous):

I also get 9; but I would let someone else try to explain before me since I'm usually incorrect

OpenStudy (anonymous):

or vice-versa

OpenStudy (anonymous):

First you do the brackets and then you go from left to right. So 6/2*3 = 9

OpenStudy (anonymous):

Jayz657 I think you did PEMDAS wrong.

OpenStudy (calculusfunctions):

It depends on how the problem is perceived. \[6\div2\times(2+1)=9\]but\[\frac{ 6 }{ 2(2+1) }=1\]

OpenStudy (anonymous):

it depends on where the (2+1) is at if its on the numerator then its 9 and if its on the denominator is 1

OpenStudy (calculusfunctions):

Exactly what I said. lol

OpenStudy (anonymous):

lol yea sorry didnt see it, was probably typing when you said that haha

OpenStudy (anonymous):

There is no fraction. The problem reads: 6 divided by 2(1+2)=? With PEMDAS I did the parenthesis (1+2)=3. Then using PEMDAS I divided 6 by 2 and got 3. Then, 3 x 3 = 9

OpenStudy (anonymous):

You are all forgetting the rule that says we go from left to right when solving equations, only if there are operations that come first, such as multiplication, division, brackets.. we switch the order. Here you do do brackets first, as always, so you get 6/2*(1+2) = 6/2*3 = (What now? You use the "Go from left to right rule.") And you get the obvious... 3*3 = 9.

OpenStudy (raden):

agree, the answer is 9

OpenStudy (anonymous):

So, my friend getting an answer of 1 is wrong because she is not using PEMDAS?

OpenStudy (anonymous):

ok w/e instead of just getting all confused with pemdas, factor in the 6/2 into the parenthesis 6/2 + 6/2(2) 3 + 6 = 9 there

OpenStudy (raden):

yea, if ur question is 6 : 2 x (3+1) really, division and multiplication as strong as of both but, first operation will be precedence it.

hartnn (hartnn):

i would go with 9

OpenStudy (anonymous):

\[6\div2\left( 1+2 \right)\]

OpenStudy (anonymous):

Thanks! I'm going with 9. My friend is sticking with 1. She's stubborn.

OpenStudy (raden):

oh, sorry... (1+2) i mean :)